# Generalization in Federated Learning: A Conditional Mutual Information Framework

### v1

@@ 1 Introduction
Recently, Yuan et al. (2022) introduces a two-level generalization framework, where client distributions are drawn from a meta-distribution. Under this framework, the generalization error is decomposed into two components: the out-of-sample gap, measuring the gap between empirical and true risk for participating clients, and the participation gap, capturing the difference in true risks between participating and non-participating clients. This broader definition extends the commonly used out-of-sample gap and is especially relevant in cross-device FL scenarios, where clients are sampled from a large, dynamic population.
Among many analytic tools, information-theoretic generalization analysis, pioneered by (Russo & Zou, 2016, 2019; Xu & Raginsky, 2017), has the advantage of accounting for both the data distribution and the learning algorithm, often giving tighter generalization bounds compared to distribution-independent or algorithm-independent methods. The original mutual information (MI) based generalization bounds in (Xu & Raginsky, 2017; Bu et al., 2020) have been applied to FL in (Yagli et al., 2020; Barnes et al., 2022; Zhang et al., 2024). However, it is well-known that these MI-based bounds suffer from unboundedness although the true generalization error can be small (Bassily et al., 2018). To address this issue, Steinke & Zakynthinou (2020) propose the conditional mutual information (CMI) framework, which introduces a “supersample” construction. Here, an auxiliary “ghost sample” is drawn alongside the original training sample, and a sequence of Bernoulli random variables determines which data points are used for training. The CMI between these Bernoulli variables and the hypothesis (e.g., model parameters), conditioned on the supersample, serves as a generalization measure. This construction ensures bounded CMI terms due to bounded entropy of Bernoulli variables, inherently leading to tighter generalization bounds compared to the standard MI-based methods, as shown in Haghifam et al. (2020). The CMI framework has seen multiple refinements, including fast-rate and numerically tight bounds (Harutyunyan et al., 2021; Hellström & Durisi, 2022a; Wang & Mao, 2023b; Dong et al., 2024). Despite the success of these CMI bounds, their effectiveness in capturing FL generalization is still not established.
@@
  • Based on our superclient and supersamples construction, We derive the first CMI-based generalization bound (Theorem 4.1) for FL, consisting of two terms: (i) the CMI between the hypothesis and the Bernoulli variable governing client participation, and (ii) the CMI between the hypothesis and the Bernoulli variable governing training data membership. These terms reveal that FL models generalize well to unseen clients when they leak minimal information about training data membership and client participation. Furthermore, we show that differential privacy constraints at both local and global levels naturally imply generalization (Lemma 4.2). In the special case of identical client data distributions, the bound reduces to a single CMI term (Corollary 4.2), aligning with intuition. We also extend our framework to high-probability bounds (Theorem 4.2) and excess risk analysis (Theorem 4.3).
  • To improve the bound’s convergence rate, we derive evaluated CMI (e-CMI) bounds (Theorem 5.1), using techniques from Wang & Mao (2023b). These bounds not only recover the best-known FL convergence rate in the low empirical risk regime, as shown in Hu et al. (2023), but also provide significant practical advantages. Unlike hypothesis-based CMI bounds that involve high-dimensional random variables, e-CMI bounds only require computing CMI between one-dimensional random variables, making them much easier to estimate in practice.
  • We extend our analysis to scenarios where model aggregation strategies (e.g., model averaging) and structured loss functions play a role. Specifically, following Barnes et al. (2022), we derive a CMI bound under Bregman loss and show that both the participation gap and out-of-sample gap exhibit fast-rate convergence with respect to the number of participating clients. Further, inspired by Gholami & Seferoglu (2024), we establish a sharper CMI bound under strongly convex and smooth loss functions, demonstrating even faster decay rates for out-of-sample gap.
@@
2 Related Work
Federated Learning Methods
FedAvg (McMahan et al., 2017) is one of the earliest and most widely used FL algorithms. However, its performance degrades significantly in non-i.i.d. settings. To mitigate this issue, FedProx (Li et al., 2020) introduces a regularization term, $||w-w_{i}||$, which helps stabilize updates and reduce divergence among local models. In fact, the squared form of this regularization term naturally arises in the derivation of our Theorem 6.2 (see Appendix F.2), and when the CMI term in Theorem 6.2 is small, this regularization term also tends to be small. SCAFFOLD (Karimireddy et al., 2020) tackles data heterogeneity using variance reduction techniques to improve convergence, while FedNova (Wang et al., 2020) introduces a normalized averaging scheme to mitigate objective inconsistencies while preserving fast training convergence. More recently, leveraging information-theoretic analysis, FedMDMI (Zhang et al., 2024) incorporates a global model-data MI term as a regularizer to improve generalization performance. Beyond these methods, numerous other FL algorithms have been proposed, we refer readers to (Huang et al., 2024; Solans et al., 2024) for recent advances.
Generalization Theory in Federated Learning
Recent research has increasingly focused on understanding generalization in FL. In addition to works in (Chen et al., 2021; Hu et al., 2023; Yagli et al., 2020; Barnes et al., 2022; Gholami & Seferoglu, 2024; Zhang et al., 2024), which we will discuss throughout the paper, several other studies provide new insights. Sefidgaran et al. (2022, 2024) introduce rate-distortion-based generalization bounds for FL, with Sefidgaran et al. (2024) demonstrating that increasing the frequency of communication rounds does not always reduce generalization error and, in some cases, may even degrade generalization performance. Stability-based generalization analysis has also been explored in FL. For instance, Sun et al. (2024) analyze the impact of data heterogeneity, concluding that greater heterogeneity leads to higher generalization error. It is also worth noting that generalization theory in meta-learning is closely related to FL. Particularly, personalized FL can be viewed as a variant of meta-learning (Fallah et al., 2020). In fact, our CMI framework is inspired in part by a recent CMI-based generalization bound for meta-learning proposed in Hellström & Durisi (2022b).
@@ 3 Preliminaries
For the $i$-th participating client, we define its true risk and empirical risk as $L_{\mu_{i}}(w)\triangleq{{\mathbb{E}_{Z\sim\mu_{i}}}\left[\ell(w,Z)\right]}$ and $L_{S_{i}}(w)=\frac{1}{n}\sum_{j=1}^{n}\ell(w,Z_{i,j})$, respectively. For the sake of simplicity, we treat each client equally, then we define the average client true risk for the participating clients as
@@
$\mathcal{E}_{\mathcal{D}}(\mathcal{A}_{[m]})\triangleq{{\mathbb{E}_{W}}\left[L_{\mathcal{D}}(W)\right]}-{{\mathbb{E}_{W,S}}\left[L_{S}(W)\right]}.$
@@
Notably the supersamples for the non-participating clients, although well defined, are actually irrelevant, as none of those data are used in training. To improve readability, Appendix A provides a visualization of our superclient and supersample construction, along with a summary of notations.
@@ 4 Hypothesis-based CMI Bound
The participation gap, denoted as $\mathcal{E}_{PG}(\mathcal{A})$, quantifies the difference in test performance between non-participating and participating clients. The out-of-sample gap, denoted as $\mathcal{E}_{OG}(\mathcal{A})$, represents the average of the local generalization gaps over the $K$ participating clients. Note that most existing FL generalization studies focus primarily on $\mathcal{E}_{OG}(\mathcal{A})$ (Yagli et al., 2020; Barnes et al., 2022; Chor et al., 2023; Sefidgaran et al., 2022, 2024; Sun et al., 2024).
@@
The participation gap $\mathcal{E}_{PG}(\mathcal{A})$ can be rewritten as,
@@
The our-of-sample gap $\mathcal{E}_{OG}(\mathcal{A})$ can be rewritten as,
@@ 4.1 First CMI Bound
We now present the first CMI bound for FL.
@@
Assume $\ell(\cdot,\cdot)\in[0,1]$, then the following bound holds
@@
Furthermore, if both CMI terms are zero, then the algorithm’s output is independent of its input (i.e., independent of the chosen participating clients and their local training data). Conversely, if they achieve their upper bounds, the algorithm reveals all membership information about the clients and their local training sets. Specifically, the first CMI term quantifies how well one can infer the “participating client set membership” from the output hypothesis when $(\tilde{z}^{i},u^{i})$ are known, whereas the second term quantifies how well one can infer the “local training set membership” when $(\tilde{z}^{i},v_{i})$ are given. Clearly, if the FL algorithm $\mathcal{A}$ enforces privacy constraints that make these inferences (i.e. determining $V_{i}$ and $U_{j}^{i,V_{i}}$) difficult, then both CMI terms remain small, resulting in a small generalization error; in other words, privacy implies generalization in FL. Inspired by (Cuff & Yu, 2016; Barnes et al., 2022), when FL algorithms are differentially private (Dwork et al., 2006a, b), this connection is formally established in the following lemma.
@@
Hence, we achieve a rate of order $\mathcal{O}\left(\frac{1}{\sqrt{K}}+\frac{1}{\sqrt{Kn}}\right)$. Recently, Zhang et al. (2024, Theorem 5.1) presents a MI-based FL bound of the form $\mathcal{O}\left(\sqrt{\frac{I(W;\mu_{[K]})}{K}}+\sqrt{\frac{I(W;S)}{Kn}}\right)$, which is the input-output mutual information (IOMI) (Xu & Raginsky, 2017) version of our Corollary 4.1. Since CMI is always no larger than the corresponding IOMI counterpart (Haghifam et al., 2020, Theorem 2.1), our bound is always tighter than Zhang et al. (2024, Theorem 5.1). Additionally, as a by-product, we also provide a novel IOMI bound for FL in Theorem C.2 in Appendix.
Before moving on, let us consider the i.i.d. (homogeneous) FL setting. Intuitively, if all clients share the same data distribution, then the first CMI term—which bounds $\mathcal{E}_{PG}(\mathcal{A})$—should vanish. The corollary below formalizes this intuition.
@@
If $|\mathcal{C}|=1$, then $\left|\mathcal{E}_{\mathcal{D}}(\mathcal{A})\right|\leq\sqrt{\frac{2I(W;U|\widetilde{Z})}{Kn}}.$
@@ 4.2 High Probability CMI Bounds
Assume $\ell(\cdot,\cdot)\in[0,1]$, let $P_{W|\widetilde{Z},U,V}$ be the the conditional distribution of $W$ given $(\widetilde{Z},U,V)$, $P_{W|\widetilde{Z},U}={{\mathbb{E}_{V}}\left[P_{W|\widetilde{Z},U,V}\right]}$ and $P_{W|\widetilde{Z},V}={{\mathbb{E}_{U}}\left[P_{W|\widetilde{Z},V}\right]}$. Then, with probability at least $1-\delta$ under the draw of $(\widetilde{Z},U,V)$, the generalization error $\left|{{\mathbb{E}_{W|\widetilde{Z},U,V}}\left[L_{\mathcal{D}}(W)\!\!-\!\!L_{S}(W))\right]}\right|$ is upper bounded by
@@
Notice that the second term in the bound has rate $\mathcal{O}(\frac{1}{\sqrt{Kn}})$ matching the PAC-Bayesian bound of Sefidgaran et al. (2024, Theorem 3.1) for bounded loss.
@@ 5 Fast-rate Evaluated CMI Bounds
The above bounds show a rate of $\mathcal{O}(\frac{1}{\sqrt{K}}+\frac{1}{\sqrt{Kn}})$, we now improve this rate. Thanks to the symmetric property of superclient and supersamples, we are able to adopt the single-loss technique from Wang & Mao (2023b). This gives us a variant of the evaluated CMI (e-CMI) bound (Steinke & Zakynthinou, 2020; Hellström & Durisi, 2022a).
@@
Notably, the square-root function is removed from the CMI terms, at the cost of a multiplicative factor $(C_{1}C_{2}-1)$ to the average client empirical risk ${{\mathbb{E}_{W,S}}\left[L_{S}(W)\right]}$. Another advantage of the e-CMI bound is that it involves only one-dimensional random variables, unlike the hypothesis-based CMI bound, where $W$ is a high-dimensional random variable. This makes e-CMI significantly easier to estimate in practice. Moreover, by the data-processing inequality, a fast-rate hypothesis-based CMI bound is established below.
@@
When the average empirical risk ${{\mathbb{E}_{W,S}}\left[L_{S}(W)\right]}$ is sufficiently small, this bound suggests a fast rate of order $\mathcal{O}\left(\frac{1}{K}+\frac{1}{Kn}\right)$. Unlike (Barnes et al., 2022; Gholami & Seferoglu, 2024), our fast-rate result does not require any specific loss function structure beyond boundedness. Similar fast-rate behavior of generalization error under additional conditions (e.g., Bernstein condition, uniform entropy on $\mathcal{W}$) is also observed by Hu et al. (2023) for Lipschitz losses. In fact, we expect these fast-rate bounds could be extended (e.g., via the variance-based and sharpness-based analyses in (Wang & Mao, 2023b; Dong et al., 2024) or binary-KL bound in (Hellström & Durisi, 2022a, b; Hellström & Guedj, 2024)), relaxing the need for strictly small empirical risk.
@@ 6 CMI Bounds for Model Aggregation in FL
While Barnes et al. (2022) proves that $\left|\mathcal{E}_{OG}(\mathcal{A})\right|$ decays with a fast decaying rate with respect to $K$, the following result shows that under a Bregman loss, the overall generalization error $\left|\mathcal{E}_{\mathcal{D}}(\mathcal{A})\right|$ for an FL algorithm using model averaging achieves such a fast convergence rate with respect to $K$, namely both $\left|\mathcal{E}_{OG}(\mathcal{A})\right|$ and $\left|\mathcal{E}_{PG}(\mathcal{A})\right|$ exhibit this fast rate.
@@
The assumptions given in (i-ii) are weaker than the boundedness assumption for the loss function, as if $\ell$ is bounded, then the difference between two losses is also bounded, which is guaranteed to be sub-Gaussian. In addition, Theorem 6.1 can be interpreted as a end-to-end generalization bound. Extending it to a multi-round setting is straightforward by following the same approach used in Barnes et al. (2022, Corollary 3).
@@
Although many popular loss functions, such as the squared loss and cross-entropy loss, can be regarded as special cases of Bregman divergence (Yamane et al., 2023), the Bregman loss $D_{f}(w,z)$ used above requires $w$ and $z$ to have the same dimension. This requirement may limit its applicability beyond nonparametric algorithms. Recently, Gholami & Seferoglu (2024) achieves an even faster decay rate for $\left|\mathcal{E}_{OG}(\mathcal{A})\right|$ by using a smooth and strongly convex loss function. Building on their assumptions and techniques, we extend this analysis in an information-theoretic framework. Additionally, we assume that each participating client’s local algorithm is an interpolating algorithm (i.e., achieving zero empirical risk). This leads to the following result.
@@
Let $\ell(\cdot,\cdot)\in[0,1]$ be $L$-smooth and $\alpha$-strongly convex in $\mathcal{W}$, and let $\mathcal{A}_{i}$ be an interpolating algorithm for each client $i$. Denote $\gamma=\frac{2L}{\alpha\log{2}}$, then
@@
Clearly, the decay rate with respect to $K$ in this bound is even faster than that for $\left|\mathcal{E}_{OG}(\mathcal{A})\right|$ in Theorem 6.1. Note that the boundedness assumption can be relax to sub-Gaussianity (i.e., an “almost bounded” case). The assumption of an interpolating algorithm is also practical due to the widespread use of overparameterized deep neural networks nowadays.
The use of smooth and strongly convex loss functions, often employed to ensure the linear convergence of gradient descent (GD), has also been studied in the context of personalized FL generalization. For example, Chen et al. (2021) shows that if the data heterogeneity—measured as the minimum average (squared) $L_{2}$ distance between global model parameters and optimal local parameters—is mild, then FedAvg achieves minimax optimality. In Theorem 6.2, data heterogeneity is captured by the factor ${\mathbb{E}\left[\ell(W,Z_{i,j})\right]}$ in the second term of the bound. In fact, if $\mathcal{A}_{i}$ does not interpolate the training data, this factor is replaced by ${\mathbb{E}\left[\ell(W,Z_{i,j})\right]}-{\mathbb{E}\left[\ell(W_{i},Z_{i,j})\right]}$. When this factor vanishes, i.e., when data heterogeneity is negligible, the second term (which contains the square-root function) becomes negligible, and the first term (with a faster rate, e.g., $\mathcal{O}\left(\frac{B}{K^{2}}\right)$ for a communication constraint) dominates the bound. This observation aligns with Chen et al. (2021), showing that FedAvg performs better when data heterogeneity is low. Additionally, we remark that extending a similar analysis to bound $\left|\mathcal{E}_{PG}(\mathcal{A})\right|$ is not straightforward. This would require that, for each client, there exists a minimizer that achieves a zero-gradient signal for any unseen local data, which is generally not guaranteed.

### v2

@@ 1 Introduction
Recently, Yuan et al. (2022) introduces a two-level generalization framework, where client distributions are drawn from a meta-distribution. Under this framework, the generalization error is decomposed into two components: the out-of-sample gap, measuring the gap between empirical and true risk for participating clients, and the participation gap, capturing the difference in true risks between participating and non-participating clients. This broader definition extends the commonly used out-of-sample gap and is especially relevant in cross-device FL scenarios, where clients are sampled from a large and dynamic population.
Among many analytic tools, information-theoretic generalization analysis, pioneered by (Russo & Zou, 2016, 2019; Xu & Raginsky, 2017), has the advantage of accounting for both the data distribution and the learning algorithm, often giving tighter generalization bounds compared to distribution-independent or algorithm-independent methods. The original mutual information (MI)-based generalization bounds in (Xu & Raginsky, 2017; Bu et al., 2020) have been applied to FL in (Yagli et al., 2020; Barnes et al., 2022; Zhang et al., 2024). However, it is well-known that these MI-based bounds suffer from unboundedness although the true generalization error can be small (Bassily et al., 2018). To address this issue, Steinke & Zakynthinou (2020) proposes the conditional mutual information (CMI) framework, which introduces a “supersample” construction. Here, an auxiliary “ghost sample” is drawn alongside the original training sample, and a sequence of Bernoulli random variables determines which data points are used for training. The CMI between these Bernoulli variables and the hypothesis (e.g., model parameters), conditioned on the supersample, serves as a generalization measure. This construction ensures bounded CMI terms due to bounded entropy of Bernoulli variables, inherently leading to tighter generalization bounds compared to the standard MI-based methods, as shown in Haghifam et al. (2020). The CMI framework has seen multiple refinements, including fast-rate and numerically tight bounds (Harutyunyan et al., 2021; Hellström & Durisi, 2022a; Wang & Mao, 2023a, c, 2024b; Dong et al., 2024). Despite the success of these CMI bounds, their effectiveness in capturing FL generalization is still not established.
@@
  • Based on our superclient and supersamples construction, We derive the first CMI-based generalization bound (cf. Theorem 3.1) for FL, consisting of two terms: (i) the CMI between the hypothesis and the Bernoulli variable governing client participation, and (ii) the CMI between the hypothesis and the Bernoulli variable governing training data membership. These terms reveal that FL models generalize well to unseen clients when they leak minimal information about training data membership and client participation. Furthermore, we show that differential privacy constraints at both local and global levels naturally imply generalization (cf. Lemma 3.2). In the special case of identical client data distributions, the bound reduces to a single CMI term (cf. Corollary 3.2), aligning with intuition. We also extend our framework to high-probability bounds (cf. Theorem 3.2) and excess risk analysis (cf. Theorem 3.3).
  • To improve the bound’s convergence rate, we derive evaluated CMI (e-CMI) bounds (cf. Theorem 4.1), using techniques from Wang & Mao (2023c). These bounds not only recover the best-known FL convergence rate in the low empirical risk regime, as shown in Hu et al. (2023), but also provide significant practical advantages. Unlike hypothesis-based CMI bounds that involve high-dimensional random variables, e-CMI bounds only require computing CMI between one-dimensional random variables, making them much easier to estimate in practice.
  • We extend our analysis to scenarios where model aggregation strategies (e.g., model averaging) and structured loss functions play a role. Specifically, following Barnes et al. (2022), we derive a CMI bound under Bregman loss and show that both the participation gap and out-of-sample gap exhibit fast-rate convergence with respect to the number of participating clients (cf. Theorem 5.1). Further, inspired by Gholami & Seferoglu (2024), we establish a sharper CMI bound under strongly convex and smooth loss functions (cf. Theorem 5.2), demonstrating even faster decay rates for out-of-sample gap.
@@ 2 Preliminaries
For the $i$-th participating client, we define its true risk and empirical risk as $L_{\mu_{i}}(w)\triangleq{{\mathbb{E}_{Z\sim\mu_{i}}}\left[\ell(w,Z)\right]}$ and $L_{S_{i}}(w)\triangleq\frac{1}{n}\sum_{j=1}^{n}\ell(w,Z_{i,j})$, respectively. For the sake of simplicity, we treat each client equally, then we define the average client true risk for the participating clients as
@@
$\mathcal{E}_{\mathcal{D}}(\mathcal{A})\triangleq{{\mathbb{E}_{W}}\left[L_{\mathcal{D}}(W)\right]}-{{\mathbb{E}_{W,S}}\left[L_{S}(W)\right]}.$
@@
Notably, the supersamples for the non-participating clients, although well defined, are actually irrelevant, as none of those data are used in training. To improve readability, Appendix A provides a visualization of our superclient and supersample construction (cf. Figure 2), along with a summary of notations (cf. Table 1).
@@ 3 Hypothesis-based CMI Bound
The participation gap, denoted as $\mathcal{E}_{PG}(\mathcal{A})$, quantifies the difference in test performance between non-participating and participating clients. The out-of-sample gap (also called participation error in (Hu et al., 2023)), denoted as $\mathcal{E}_{OG}(\mathcal{A})$, represents the average of the local generalization gaps over the $K$ participating clients. Note that most existing FL generalization studies focus primarily on $\mathcal{E}_{OG}(\mathcal{A})$ (Yagli et al., 2020; Barnes et al., 2022; Chor et al., 2023; Sefidgaran et al., 2022, 2024; Sun et al., 2024).
@@
The participation gap $\mathcal{E}_{PG}(\mathcal{A})$ can be rewritten as
@@
The our-of-sample gap $\mathcal{E}_{OG}(\mathcal{A})$ can be rewritten as
@@ 3.1 First CMI Bound
The bounding steps for $\mathcal{E}_{\mathcal{D}}(\mathcal{A})$ closely follow those in standard centralized analysis, once Lemma 3.1 has been established.
We are now in a position to present the first CMI bound for FL.
@@
Assume $\ell(\cdot,\cdot)\in[0,1]$, the following bound holds
@@
Furthermore, if both CMI terms are zero, then the algorithm’s output is independent of its input (i.e., independent of the chosen participating clients and their local training data). In contrast, if they achieve their upper bounds, the algorithm reveals all membership information about the clients and their local training sets. Specifically, the first CMI term quantifies how well one can infer the membership of the “participating client set” from the output hypothesis when $(\tilde{z}^{i},u^{i})$ are known, whereas the second term quantifies how well one can infer the membership of the “local training set” when $(\tilde{z}^{i},v_{i})$ are given. Clearly, if the FL algorithm $\mathcal{A}$ enforces privacy constraints that make these inferences (i.e. determining $V_{i}$ and $U_{j}^{i,V_{i}}$) difficult, then both CMI terms remain small, resulting in a small generalization error; in other words, privacy implies generalization in FL. Inspired by (Cuff & Yu, 2016; Barnes et al., 2022), when FL algorithms are differentially private (Dwork et al., 2006a, b), this connection is formally established in the following lemma.
@@
Hence, we achieve a rate of order [^1] In this paper, due to the absence of explicit decay rates for the CMI or MI terms, we often follow previous works by simply treating these terms as $\mathcal{O}(1)$ when stating order-wise behavior. However, it should be noted that the actual decay of the bound is likely to be faster than the rate presented. $\mathcal{O}\left(\frac{1}{\sqrt{K}}+\frac{1}{\sqrt{Kn}}\right)$. Recently, Zhang et al. (2024, Theorem 5.1) presents a MI-based FL bound of the form $\mathcal{O}\left(\sqrt{\frac{I(W;\mu_{[K]})}{K}}+\sqrt{\frac{I(W;S)}{Kn}}\right)$, which is the input-output mutual information (IOMI) (Xu & Raginsky, 2017) version of our Corollary 3.1. Since CMI is always no larger than the corresponding IOMI counterpart (Haghifam et al., 2020, Theorem 2.1), our bound is always tighter than Zhang et al. (2024, Theorem 5.1). Additionally, as a by-product, we also provide a novel IOMI bound for FL in Theorem C.2 in Appendix.
Before moving on, let us consider the i.i.d. (homogeneous) FL setting. Intuitively, if all clients share the same data distribution, then the first CMI term—which bounds $\mathcal{E}_{PG}(\mathcal{A})$—should vanish. In particular, when each $\mu_{i}$ is identical, $V_{i}$ no longer influences the algorithm’s output distribution once $\widetilde{Z}^{i}$ and $U^{i}$ are given. The corollary below formalizes this intuition.
@@
If $|\mathcal{C}|=1$, then we have
$\left|\mathcal{E}_{\mathcal{D}}(\mathcal{A})\right|\leq\sqrt{\frac{2I(W;U|\widetilde{Z})}{Kn}}.$
@@ 3.2 High Probability CMI Bounds
Assume $\ell(\cdot,\cdot)\in[0,1]$, and let $P_{W|\widetilde{Z},U,V}$ be the the conditional distribution of $W$ given $(\widetilde{Z},U,V)$. Let $P_{W|\widetilde{Z},U}={{\mathbb{E}_{V}}\left[P_{W|\widetilde{Z},U,V}\right]}$ and let $P_{W|\widetilde{Z},V}={{\mathbb{E}_{U}}\left[P_{W|\widetilde{Z},V,U}\right]}$. Then, with probability at least $1-\delta$ under the draw of $(\widetilde{Z},U,V)$, the generalization error $\left|{{\mathbb{E}_{W|\widetilde{Z},U,V}}\left[L_{\mathcal{D}}(W)\!\!-\!\!L_{S}(W))\right]}\right|$ is upper bounded by
@@
Notice that the second term in the bound has the rate $\mathcal{O}(\frac{1}{\sqrt{Kn}})$ matching the PAC-Bayesian bound of Sefidgaran et al. (2024, Theorem 3.1) for a bounded loss.
@@ 4 Fast-rate Evaluated CMI Bounds
The above bounds show a rate of $\mathcal{O}(\frac{1}{\sqrt{K}}+\frac{1}{\sqrt{Kn}})$, we now improve this rate. Thanks to the symmetric property of superclient and supersamples, we are able to adopt the single-loss technique from Wang & Mao (2023c). This gives us a variant of the evaluated CMI (e-CMI) bound (Steinke & Zakynthinou, 2020; Hellström & Durisi, 2022a).
@@
Notably, the square-root functions are removed from the CMI terms, at the cost of a multiplicative factor $C_{1}C_{2}$ to the average client empirical risk ${{\mathbb{E}_{W,S}}\left[L_{S}(W)\right]}$. Another advantage of the e-CMI bound is that it involves only one-dimensional random variables, unlike the hypothesis-based CMI bound, where $W$ is a high-dimensional random variable. This makes e-CMI significantly easier to estimate in practice. Moreover, by the data-processing inequality, a fast-rate hypothesis-based CMI bound is established below.
@@
When the average empirical risk ${{\mathbb{E}_{W,S}}\left[L_{S}(W)\right]}$ is sufficiently small, this bound suggests a fast rate of order $\mathcal{O}\left(\frac{1}{K}+\frac{1}{Kn}\right)$. Unlike (Barnes et al., 2022; Gholami & Seferoglu, 2024), our fast-rate result does not require any specific loss function structure beyond boundedness. Similar fast-rate behavior of generalization error under additional conditions (e.g., Bernstein condition, uniform entropy on $\mathcal{W}$) is also observed by Hu et al. (2023) for Lipschitz losses. In fact, we expect our fast-rate bounds could be extended (e.g., via the variance-based and sharpness-based analyses in (Wang & Mao, 2023c; Dong et al., 2024) or binary-KL bound in (Hellström & Durisi, 2022a, b; Hellström & Guedj, 2024)), relaxing the need for strictly small empirical risk.
@@ 5 CMI Bounds for Model Aggregation in FL
While Barnes et al. (2022) proves that $\left|\mathcal{E}_{OG}(\mathcal{A})\right|$ decays with a fast rate with respect to $K$, the following result shows that under a Bregman loss, the overall generalization error $\left|\mathcal{E}_{\mathcal{D}}(\mathcal{A})\right|$ for an FL algorithm using model averaging achieves such a fast convergence rate with respect to $K$, namely both $\left|\mathcal{E}_{OG}(\mathcal{A})\right|$ and $\left|\mathcal{E}_{PG}(\mathcal{A})\right|$ exhibit this fast rate.
@@
The assumptions given in (i-ii) are weaker than the boundedness assumption for the loss function, as if $\ell$ is bounded, then the difference between two losses is also bounded, which is guaranteed to be sub-Gaussian. In the case of heavy-tailed losses, additional techniques such as truncation (Dong et al., 2024) would be required to ensure the results hold. Furthermore, Theorem 5.1 can be interpreted as an end-to-end generalization bound for a single round of FedAvg. Extending it to a multi-round setting is straightforward by following the same approach used in Barnes et al. (2022, Corollary 3).
@@
Although many popular loss functions, such as the squared loss and KL divergence, can be regarded as special cases of Bregman divergence (Yamane et al., 2023), the Bregman loss $D_{f}(w,z)$ used above requires $w$ and $z$ to have the same dimension. This requirement may limit its applicability beyond nonparametric algorithms. Recently, Gholami & Seferoglu (2024) achieves an even faster decay rate for $\left|\mathcal{E}_{OG}(\mathcal{A})\right|$ by using a smooth and strongly convex loss function. Building on their assumptions and techniques, we extend this analysis in an information-theoretic framework. Additionally, we assume that each participating client’s local algorithm is an interpolating algorithm (i.e., achieving zero empirical risk). This leads to the following result.
@@
Let $\ell(\cdot,\cdot)\in[0,1]$ be $L$-smooth and $\alpha$-strongly convex in $\mathcal{W}$, and let $\mathcal{A}_{i}$ be an interpolating algorithm for each client $i$. Let $\gamma=\frac{2L}{\alpha\log{2}}$, then
@@
Clearly, the decay rate with respect to $K$ in this bound is even faster than that for $\left|\mathcal{E}_{OG}(\mathcal{A})\right|$ in Theorem 5.1. Note that the boundedness assumption can be relax to sub-Gaussianity (i.e., an “almost bounded” case). The assumption of an interpolating algorithm is also practical due to the widespread use of overparameterized deep neural networks in modern deep learning scenarios.
Regarding the local CMI term $I(W_{i};U_{j}^{i,V_{i}}|\widetilde{Z}^{i},V_{i})$ in Theorem 5.2, note that in our single-round setting, the local model is trained independently and does not communicate with other clients. As a result, the CMI term associated with client $i$ is, by construction, independent of the CMI terms of other clients. It reflects only the relationship between a single client’s data and its corresponding model and therefore does not, in general, provide a direct or definitive measure of cross-client data heterogeneity. However, if the data distribution itself is governed by a heterogeneity parameter, as in the experimental setup of Sun et al. (2024), then the value of this CMI term may indeed vary with that parameter, as it ultimately depends on the underlying data distribution.
Moreover, the use of smooth and strongly convex loss functions, often employed to ensure the linear convergence of gradient descent (GD), has also been studied in the context of personalized FL generalization. For example, Chen et al. (2021) shows that if the data heterogeneity—measured as the minimum average (squared) $L_{2}$ distance between global model parameters and optimal local parameters—is mild, then FedAvg achieves minimax optimality. In Theorem 5.2, data heterogeneity is captured by the factor ${\mathbb{E}\left[\ell(W,Z_{i,j})\right]}$ in the second term of the bound. In fact, if $\mathcal{A}_{i}$ does not interpolate the training data, this factor is replaced by ${\mathbb{E}\left[\ell(W,Z_{i,j})\right]}-{\mathbb{E}\left[\ell(W_{i},Z_{i,j})\right]}$. When this factor vanishes, i.e., when data heterogeneity is negligible, the second term (which contains the square-root function) becomes negligible, and the first term (with a faster rate, e.g., $\mathcal{O}\left(\frac{B}{K^{2}}\right)$ for a communication constraint) dominates the bound. This observation aligns with Chen et al. (2021), showing that FedAvg performs better when data heterogeneity is low.
Additionally, we remark that extending a similar analysis to bound $\left|\mathcal{E}_{PG}(\mathcal{A})\right|$ is not straightforward. This would require that, for each client, there exists a minimizer that achieves a zero-gradient signal for any unseen local data, which is generally not guaranteed.